Exercise


May 08'23

Answer

Solution: D

If a policy is of Type A, the probability that the two claims are equal is

(0.4)(0.4) + (0.3)(0.3) + (0.2)(0.2) + (0.1)(0.1) = 0.16 + 0.09 + 0.04 + 0.01 = 0.30.

If a policy is of Type B, the probability that the two claims are equal is 4(0.25)(0.25) = 0.25. Therefore, the probability that a randomly selected policy has equal claims is

0.70(0.30) + 0.30(0.25) = 0.285.

If four policies are selected, the desired probability is the probability that a binomial random variable with n = 4 and p = 0.285 is 1. This is

4(0.285)(1 – 0.285)3 = 0.417.

Copyright 2023. The Society of Actuaries, Schaumburg, Illinois. Reproduced with permission.

00
Comments
You are not permitted to add comments. Make sure you are logged in and your email has been confirmed.